what is the probability that of 2 randomly selected traffic fatalities in 2009 at least one was alcohol related
Answer #1

P(alcohol related) = 9820/33800 = 0.29

P(not alcohol related) = 1 - 0.29 = 0.71

P(both are alcohol related)=0.29 * 0.29 = 0.084

P(neither of the two are alcohol related) = 0.71 * 0.71 = 0.504

Ans:

P(at least one was alcohol related) = 0.29 * 0.71 + 0.71 * 0.29 + 0.29 * 0.29 = 0.496